The commodity and money markets of Juja farm, a hypothetical economy are as



given below:




Commodity Market



Y = C + I



C = 1000 + 0.8Y



I = 2000 – 0.75r



Money Market



Lt = 0.25Y (Transactions and precautionary demand for money function)



Ls = 1000 – 0.5r (Speculative demand for money function)



Ms = 3200 (Money supply function)



Required:



i) Mathematically derive both IS and LM curves. (6 marks)



ii) Derive the equilibrium level of Income and rate of interest.



(4 marks)



iii) If the money supply is increased by 80, what would be the effect on the equilibrium level of income and rate of interest in Juja farm economy?




1
Expert's answer
2022-03-28T14:06:04-0400

i)

IS= Y = C + I

Y=1000 + 0.8Y+2000 – 0.75r

0.2Y= 3000-0.75r

Y= 15000-3.75r

Ms= Md=

0.25Y+1000 – 0.5r=3200

0.25Y=2200+0.5r

Y= 8800+2r

ii)LM=IS

8800+2r= 15000-3.75r

5.75r= 6200

r= 1,078.26

Y= 8800+2(1078.26)

Y= 10,956.52

iii) If money supply increased by 80

the new Money supply= 3280

"\therefore" 0.25Y+1000 – 0.5r=3280

0.25Y= 2280+ 0.5r

Y= 9120+2r

LM=IS

9120+2r= 15000-3.75r

5.75r= 5880

r= 1,022.61

Y= 9120+2(1022.61)

Y= 11,165.22